OpenStudy (anonymous):

(x^9-3)+(X^9+3)=X^81-9?

OpenStudy (anonymous):

Is it multiplication or addition?

OpenStudy (ash2326):

tena is that + or * in between the two terms

OpenStudy (anonymous):

x^9-3 + x^9 + 3 = 2x^9 But, (x^9-3)(x^9+3) = x^18 - 9

OpenStudy (anonymous):

well i had x^18 then i changed it lol

OpenStudy (anonymous):

2x^2-9?

OpenStudy (anonymous):

Where do you get 2x^2?

OpenStudy (anonymous):

oh sorry i ment 2x^9